A+Bi
(-5+2i)+(14-3i)

Answer by jim_thompson5910(35256)

(-5+2i)+(14-3i)


-5+2i+14-3i


(-5+14)+(2i-3i)


9-i


So (-5+2i)+(14-3i) = 9-i


So the expression is now in A+Bi form where A = 9 and B = -1


If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim